### v2

@@ 2 Related work
This accelerated matrix-vector multiplication becomes the basis for solving linear systems using iterative techniques such as GMRES [33], CG [25, 36], BiCGstab [40], and other iterative solvers. Iterative methods, while versatile, have a drawback: their convergence rate depends on the conditioning or eigenvalue clustering of the matrix. Thus, preconditioning for iterative methods, where the matrix is in $\mathcal{H}^{2}$matrix format, is often a necessity [43, 45]. In contrast, fast direct solvers guarantee a predefined number of operations for solving the system. Although the complexity for general dense matrices remains $\mathcal{O}(N^{3})$, leveraging the hierarchical matrix format has led to breakthroughs in direct solver efficiency. Researchers in [22] introduced an $\mathcal{O}(N\log N)$ direct solver algorithm, albeit with a substantial constant. Subsequent works [2, 31, 37] utilizing $\mathcal{H}^{2}$ (FMM) format have achieved $\mathcal{O}(N)$ direct solver algorithms with more favorable constants.
@@
Multigrid (MG) methods, as well as rank structured methods, play a crucial role in solving large linear systems, particularly systems with sparse matrices arising from discretized partial differential equations. Multigrid has its origins in [15] and was further developed in [9]. Multigrid has become a cornerstone of numerical techniques for solving large sparse linear systems. Its strategy lies in efficiently reducing errors at multiple scales by leveraging a hierarchy of coarser grids. Over time, a variety of fruitful generalizations have emerged, such as algebraic multigrid (AMG) [10] approaches that extend its applicability to unstructured grids and irregular geometries [14, 12, 19] and fast multipole preconditioners for sparse matrices [26]. Multigrid methods have also been turned into $\mathcal{H}^{2}$ matrices in order to provide fast methods for evaluating integral operators [5].
@@ 3.1 $\mathcal{H}^{2}$ matrix
$[D]_{ij}=\begin{cases}\widehat{D}_{ij},&\quad\text{if }A_{ij}\text{ is a closeblock}\\ 0,&\quad\text{if }A_{ij}\text{ is a far block}\end{cases}.$
@@
$[F]_{ij}=\begin{cases}0,&\quad\text{if }A_{ij}\text{ is a close block}\\ \widehat{F}_{ij}=\widehat{U}_{i}\widehat{S}_{ij}\widehat{V}_{j},&\quad\text{if}A_{ij}\text{ is a far block}\end{cases}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}[S_{1}]_{ij}}=\begin{cases}0,&\quad\text{if }A_{ij}\text{ is a close block}\\ \widehat{S}_{ij},&\quad\text{if }A_{ij}\text{ is a far block}\end{cases}.$
@@
The matrix $S_{1}$ is a dense matrix. Our next goal is to compress it using the same strategy as we applied to matrix $A$. We assemble together blocks of matrix $S_{1}$ in patches of size $p$. (Typical choices of $p$ are $2$ or $2^{d}$, where $d$ is the physical dimension of the problem.) We obtain matrix $S_{1}$ with $\frac{M}{p}$ blocks. Those blocks again separate into a set of full-rank close and low-rank far blocks. Analogously to the previous procedure, we obtain the factorization
@@
Figure 3: Schematic of $\mathcal{H}^{2}$ matrix-by vector multiplication, $l=2$
Remark 1.
@@ 3.3 $\mathcal{H}^{2}$-MG
Remark 2.
Note that $A_{2}\in\mathbb{R}^{N_{2}\times N_{2}}$ has $\mathcal{H}^{2}$ structure, just like the matrix $A$. The matrices $A$ and $A_{2}$ have exactly the same structure, except that the matrix $A$ stores the full matrix $D$, while $A_{2}$ stores a reduced part of it, $U_{1}^{\top}DV^{\top}$. Thus, matrix $A_{2}$ stores less information than $A$. Since $A$ scales linearly as an $\mathcal{H}^{2}$ matrix, $A_{2}$ also scales linearly. See the detailed proof in Section 3.4.
@@
$e_{l-1}=\textbf{Iter}(A_{l-1},r_{l-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tilde{e}_{l-1}}).$
@@ 3.4 Complexity analysis
The crucial feature of the $\mathcal{H}^{2}$-MG complexity analysis is the linear complexity of the $\mathcal{H}^{2}$ matrix. According to Remark 1, for the $\mathcal{H}^{2}$ matrix of size $N\times N$, the construction complexity, the memory requirements, and matrix-by-vector product complexity are all $\mathcal{O}(N)$. Assume $c_{H2}$ to be the $\mathcal{H}^{2}$ matrix-vector product constant.
@@ 4 Numerical results
In the kernel matrices tests, we randomly generate $x_{\text{true}}$, then compute $b=Ax_{\text{true}}$ and solve the system $Ax=b$ for $x$. Then we plot convergence of the $A$-norm of the error $||e_{k}||_{A}$ vs. iteration count $k$, where the $A$-norm refers to the energy norm induced by matrix $A$.
Remark 5.
@@ 4.1 Gaussian kernel
where $\sigma\in\mathbb{R}$ is the dispersion parameter, $c\in\mathbb{R}$ is a small regularization parameter. Matrix $A$ is approximated to the $\mathcal{H}^{2}$ format with accuracy $\epsilon=10^{-9}$ ($\epsilon$ is a relative error between the result of a matrix-vector multiplication performed using the $\mathcal{H}^{2}$ matrix and that using the original matrix), and the number of levels is chosen adaptively.
@@ 4.1.1 Gaussian matrix, analysis of error evolution
To analyze the components of the error after smoothing in the basis of the interpolation matrix $U_{1}$, or in other words, the part of the error that will be projected to the coarser levels, we compute $\text{pinv}(U_{1})e_{1*}$. Similarly, for the part of the after-smoothing error in the basis of the matrix $Q_{1}$, $\text{pinv}(Q_{1})e_{1e}$.
We study the 1D Gaussian matrix with points equally spaced between 00 and $1$, $N=1024$, $\sigma=0.01$, $c=10^{-3}$. The results of this analysis are presented in Figure 5. The blue curves represent the error components after the first smoothing step in the initial $V$-cycle, the red curves are for the second V-cycle.
(a) Component of the post-smoothing error in the basis of $U_{1}$
(b) Component of the post-smoothing error in the basis of $Q_{1}$
@@
The graphs clearly shows that the error component $\text{pinv}(U_{1})e_{1*}$ is larger than that in the error component $\text{pinv}(Q_{1})e_{1*}$ components, demonstrating that the CG smoothing can complement coarse grid correction effectively.
@@ 4.1.2 Gaussian kernel matrix, analysis of number of smoothing steps
Consider the system (9) with matrix $A$ given by (10) with $c=10^{-3}$, $\sigma=0.1$. In Remark 4, we considered assigning the finest level of the problem a different number of smoothing steps than the coarser levels. In this subsection, we analyze the effect of parameters $n_{f}$ and $n_{c}$ on the $\mathcal{H}^{2}$-MG convergence.
We first study the convergence rate of the $\mathcal{H}^{2}$-MG depending on the number of fine iterations. Figure 6 presents the convergence comparison for different numbers of fine iterations, $n_{f}$ to the accuracy $\varepsilon=10^{-9}$, where $\varepsilon=\frac{(eAe^{\top})^{\frac{1}{2}}}{||b||_{2}}<10^{-9}$. The number of coarse iterations is fixed $n_{c}=40$. Fine and coarse iteration comparisons are given for matrix size $N=8\times 10^{4}$. For $\mathcal{H}^{2}$-MG , we can track only the $A$-norm of the error on the finest level, which makes it especially useful for comparison with CG. In all figures below, we plot this error per finest level iterations (outer iterations). Since coarse-level iterations are not shown directly in these plots, we also include error-per-time plots to reflect the computational effort spent on coarser levels.
(a) $A$-norm of the error per coarse grid iteration
(b) $A$-norm of the error per time
@@
We can see that in terms of convergence per iteration, fewer fine grid iterations lead to better convergence. Figure 7 presents the convergence comparison to the accuracy $\varepsilon=10^{-9}$ for different numbers of coarse iterations, $n_{c}$. The number of fine iterations is fixed at $n_{f}=1$.
(a) $A$-norm of the error per coarse grid iteration
(b) $A$-norm of the error per time
@@
In this case, the number of coarse iterations directly influences the speed of convergence: the more coarse iterations, the faster the convergence. However, in terms of time, the optimal number of coarse iterations is an intermediate value ($n_{c}=40$). This is because there is a trade-off between the time spent on additional coarse iterations and the resulting improvement in convergence speed.
Remark 6.
@@ 4.3 Boundary element method for a 3D electrostatic problem
[…]

### v3

@@ 2 Related work
This accelerated matrix-vector multiplication becomes the basis for solving linear systems using iterative techniques such as GMRES [36], CG [26, 39], BiCGstab [43], and other iterative solvers. Iterative methods, while versatile, have a drawback: their convergence rate depends on the conditioning or eigenvalue clustering of the matrix. Thus, preconditioning for iterative methods, where the matrix is in $\mathcal{H}^{2}$ -matrix format, is often a necessity [46, 49]. In contrast, fast direct solvers guarantee a predefined number of operations for solving the system. Although the complexity for general dense matrices remains $\mathcal{O}(N^{3})$, leveraging the hierarchical matrix format has led to breakthroughs in direct solver efficiency. Researchers in [23] introduced an $\mathcal{O}(N\log N)$ direct solver algorithm, albeit with a substantial constant. Subsequent works [2, 34, 40, 30, 31, 7, 47] utilizing $\mathcal{H}^{2}$ (FMM) format have achieved $\mathcal{O}(N)$ direct solver algorithms with more favorable constants.
@@
Multigrid (MG) [10, 21] methods, as well as rank structured methods, play a crucial role in solving large linear systems, particularly systems with sparse matrices arising from discretized partial differential equations. Multigrid has its origins in [16] and was further developed in [10]. Multigrid has become a cornerstone of numerical techniques for solving large sparse linear systems. Its strategy lies in efficiently reducing errors at multiple scales by leveraging a hierarchy of coarser grids. Over time, a variety of fruitful generalizations have emerged, such as algebraic multigrid (AMG) [11] approaches that extend its applicability to unstructured grids and irregular geometries [15, 13, 20] and fast multipole preconditioners for sparse matrices [27]. Multigrid methods have also been turned into $\mathcal{H}^{2}$ matrices in order to provide fast methods for evaluating integral operators [5].
@@ 3.1 $\mathcal{H}^{2}$ matrix
[…]
@@
[…]
@@
[…]
@@
The matrix $S_{1}$ is a dense matrix. Our next goal is to compress it using the same strategy as we applied to matrix $A$. We assemble together blocks of matrix $S_{1}$ in patches of size $p$. (Typical choices of $p$ are $2$ or $2^{d}$, where $d$ is the physical dimension of the problem.) We obtain matrix $S_{1}$ with $\frac{M}{p}$ block rows and $\frac{M}{p}$ block columns. Those blocks again separate into a set of full-rank close and low-rank far blocks. Analogously to the previous procedure, we obtain the factorization
@@
Figure 3: Schematic of $\mathcal{H}^{2}$ matrix-by-vector multiplication, $l=2$
Remark 3.1.
@@ 3.3 $\mathcal{H}^{2}$-MG
Remark 3.2.
Note that $A_{2}\in\mathbb{R}^{N_{2}\times N_{2}}$ has $\mathcal{H}^{2}$ structure, just like the matrix $A$. The matrices $A$ and $A_{2}$ have exactly the same structure, except that the matrix $A$ stores the full matrix $D$, while $A_{2}$ stores a reduced part of it, $U_{1}^{\top}DV_{1}^{\top}$. Thus, matrix $A_{2}$ stores less information than $A$. Since $A$ scales linearly as an $\mathcal{H}^{2}$ matrix, $A_{2}$ also scales linearly. See the detailed proof in Section 3.4.
@@
$e_{l-1}=\textbf{Iter}(A_{l-1},r_{l-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{e}_{l-1}}).$
@@
2:for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}=\texttt{1\ldots(}l\texttt{-1)}$ do
@@
7:for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}\texttt{ = (}l\texttt{-1)\ldots 1}$ do
@@ 3.4 Complexity analysis
The crucial feature of the $\mathcal{H}^{2}$-MG complexity analysis is the linear complexity of the $\mathcal{H}^{2}$ matrix. According to Remark 3.1, for the $\mathcal{H}^{2}$ matrix of size $N\times N$, the construction complexity, the memory requirements, and matrix-by-vector product complexity are all $\mathcal{O}(N)$. Assume $c_{H2}$ to be the $\mathcal{H}^{2}$ matrix-vector product constant.
@@ 4 Numerical results
In the kernel matrix tests, we randomly generate $x_{\text{true}}$, then compute $b=Ax_{\text{true}}$ and solve the system $Ax=b$ for $x$. Then we plot convergence of the $A$-norm of the error $e_{k}||_{A}$ vs. iteration count $k$, where the $A$-norm refers to the energy norm induced by matrix $A$.
Remark 4.1.
@@ 4.1 Gaussian kernel
where $\sigma\in\mathbb{R}$ is the dispersion parameter, $c\in\mathbb{R}$ is a small regularization parameter. Matrix $A$ is approximated in the $\mathcal{H}^{2}$ format with accuracy $\epsilon=10^{-9}$ ($\epsilon$ is a relative error between the result of a matrix-vector multiplication performed using the $\mathcal{H}^{2}$ matrix and that using the original matrix), and the number of levels is chosen adaptively.
@@ 4.1.1 Gaussian matrix, analysis of error evolution
To analyze the components of the error after smoothing in the basis of the interpolation matrix $U_{1}$, or in other words, the part of the error that will be projected to the coarser levels, we compute $(U_{1})^{\top}e_{1*}$. Similarly, for the part of the after-smoothing error in the basis of the matrix $Q_{1}$, $(Q_{1})^{\top}e_{1*}$.
We study the 1D Gaussian matrix with points equally spaced between 0 and $1$, $N=1024$, $\sigma=0.01$, $c=10^{-3}$. The results of this analysis are presented in Figure 5. The blue curves represent the error components after the first smoothing step in the initial $V$-cycle, the red curves are for the second V-cycle.
(a) Vector components of $(U_{1})^{\top}e_{1*}$
(post-smoothing error in the basis of $U_{1}$)
(b) Vector components of $(Q_{1})^{\top}e_{1*}$
(post-smoothing error in the basis of $Q_{1}$)
@@
The graphs clearly shows that the error component $(U_{1})^{\top}e_{1*}$ is larger than that in the error component $(Q_{1})^{\top}e_{1*}$ components, demonstrating that the CG smoothing can complement coarse grid correction effectively.
@@ 4.1.2 Gaussian kernel matrix, analysis of number of smoothing steps
Consider the system (9) with matrix $A$ given by (10) with $c=10^{-3}$, $\sigma=0.1$. In Remark 3.4, we considered assigning the finest level of the problem a different number of smoothing steps than the coarser levels. In this subsection, we analyze the effect of parameters $n_{f}$ and $n_{c}$ on the $\mathcal{H}^{2}$-MG convergence.
[…]
(a) AA-norm of the error per coarse grid iteration
(b) AA-norm of the error per time
@@
We can see that using a smaller $n_{f}$ parameter leads to faster $A$-norm convergence, both when measured per outer iteration and per time.
Figure 7 presents the convergence comparison to the accuracy $\varepsilon=10^{-9}$ for different numbers of coarse iterations, $n_{c}$. The number of fine iterations is fixed at $n_{f}=1$.
(a) AA-norm of the error per coarse grid iteration
(b) AA-norm of the error per time
@@
In this case, the $n_{c}$ parameter directly influences the speed of $A$-norm convergence per outer iteration: the more $n_{c}$, the faster the convergence per outer iteration. However, in terms of convergence per time, the optimal number of coarse iterations is an intermediate value ($n_{c}=40$). This is because there is a trade-off between the time spent on additional coarse iterations and the resulting improvement in convergence speed.
Remark 4.2.
@@ 4.3 Boundary element method for a 3D electrostatic problem
$u(\mathbf{x})=\int_{\Gamma}\frac{1}{4\pi\|\mathbf{x}-\mathbf{y}\|}\,\sigma(\mathbf{y})\,dS_{\mathbf{y}},$